# A user-friendly package and workflow for generating effective homogeneous rheologies for the study of the long-term orbital evolution of multilayered planetary bodies

### 29 Mar 2026 v2

@@ 4.1 The homogeneous model
Table 4: Secular relaxation modes extracted for the 5-layer lunar model, sorted by relaxation timescale.
@@
Table 5: Extracted parameters for the Kelvin–Voigt elements of the full effective lunar model. Elements retained in the reduced model are marked with an asterisk.
@@ 4.2 Dominant-mode reduction over a frequency interval
(a) Real part of the Love number, $\operatorname{Re}[k_{2}(\omega)]$.
(b) Imaginary part of the Love number, $\operatorname{Im}[k_{2}(\omega)]$.
@@
(a) Real part of the Love number, $\operatorname{Re}[k_{2}(\omega)]$.
(b) Imaginary part of the Love number, $\operatorname{Im}[k_{2}(\omega)]$.
@@ 5 Conclusions
We have presented a Wolfram Language package that automates the construction of effective homogeneous generalized Voigt rheologies for incompressible, spherically symmetric layered viscoelastic bodies. The package combines three components: the forward computation of the tidal Love number for a multilayer model, the numerical extraction of its secular poles and residues, and the inversion of the layered response into equivalent homogeneous rheology parameters. Starting from a user-defined layered model, the package returns the parameters of the corresponding homogeneous generalized Voigt body in a format suitable for validation, comparison, and downstream numerical use and applications.
The lunar case study illustrates the purpose of the software: to transform an interior model into a reproducible and computationally efficient homogeneous rheology. For a five-layer model including both solid and liquid regions, it recovers the equivalent homogeneous rheology together with a reduced model obtained through dominant-mode selection over a prescribed frequency interval. This reduction is particularly useful when the relevant forcing frequencies lie in a limited range, since it preserves the tidal response of the layered model while reducing the number of rheological elements.
The package allows detailed interior-structure models to be translated into generalized Voigt parameters that can be supplied directly to time-domain tidal evolution codes, enabling their dissipation properties to be incorporated in long-term orbital evolution calculations without repeatedly solving the full layered deformation problem during integration. We expect it to be useful for studies of the Moon, terrestrial planets, and icy satellites, especially in contexts where tidal dissipation must be coupled to orbital or rotational evolution over long timescales. Future developments may include extensions to broader classes of rheologies and tighter integration with external dynamical solvers.
@@ Acknowledgements.
The author would like to thank Clodoaldo Grotta Ragazzo (USP) for many discussions and long-standing collaboration on this topic, Diogo Aguiar Gomes (KAUST) for his kind support, and J. Ricardo G. Mendonça (USP) for technical advice and a careful reading of the manuscript that greatly improved its presentation. This work was partially supported by the King Abdullah University of Science and Technology (KAUST), Saudi Arabia.

### 03 May 2026 v3

@@ 4.1 The homogeneous model
Table 4: Secular relaxation modes extracted for the 5-layer lunar model. [^5] Modes are sorted by relaxation timescale.
@@
Table 5: Extracted parameters for the Kelvin–Voigt elements of the full effective lunar model. [^6] Elements retained in the reduced model are marked with an asterisk.
@@ 4.2 Dominant-mode reduction over a frequency interval
(a) Real part of the Love number, $\operatorname{Re}[k_{2}(i\omega)]$.
(b) Imaginary part of the Love number, $\operatorname{Im}[k_{2}(i\omega)]$.
@@
(a) Real part of the Love number, $\operatorname{Re}[k_{2}(i\omega)]$.
(b) Imaginary part of the Love number, $\operatorname{Im}[k_{2}(i\omega)]$.
@@ 5 Conclusions
We have presented a Wolfram Language package that constructs effective homogeneous generalized Voigt rheologies for incompressible, spherically symmetric layered viscoelastic bodies. The package computes the tidal Love number of a multilayer model, extracts its secular poles and residues, and converts this information into the parameters of an equivalent homogeneous rheology. Starting from a user-defined layered model, the package returns the parameters of the corresponding homogeneous generalized Voigt body in a format suitable for validation, comparison, and downstream numerical use and applications.
The lunar case study shows how the code converts a five-layer interior model into an equivalent homogeneous rheology. When all extracted Voigt elements are maintained, the homogeneous model reproduces the layered tidal response within numerical precision. The dominant-mode selection then gives a smaller model adapted to the prescribed frequency interval. The reduction applies when the relevant forcing frequencies lie in a limited range, since it preserves the tidal response of the layered model while reducing the number of rheological elements.
The package allows detailed interior-structure models to be translated into generalized Voigt parameters that can be supplied directly to time-domain tidal evolution codes, enabling their dissipation properties to be incorporated in long-term orbital evolution calculations without repeatedly solving the full layered deformation problem during integration. The same procedure can be applied to other incompressible multilayer models, including models of terrestrial planets and icy satellites, provided that the layer rheologies satisfy the assumptions used in the reduction.
The current implementation assumes a Maxwell rheology for each solid layer. Maxwell bodies underestimate tidal dissipation at forcing frequencies close to or above the inverse Maxwell time, where laboratory and seismological evidence favors rheologies with a broader distribution of relaxation times, such as Andrade or Sundberg–Cooper. Our reduction pipeline is exact for a finite multilayer Maxwell body because the resulting $k_{2}(s)$ is rational with a finite set of secular poles, a property that rheologies with a continuous relaxation spectrum do not share. Andrade and similar rheologies can nevertheless be approximated within a prescribed frequency range by a finite set of dominant modes, as done in Gevorgyan et al. (2020) for an extended Burgers model. Extending the package in this direction, together with tighter integration with external dynamical solvers, will require extending the current pole–residue reduction beyond the finite Maxwell case.
@@ Acknowledgements.
The author would like to thank the anonymous referee for careful reading and constructive comments that improved the manuscript. The author would also like to thank Clodoaldo Ragazzo (USP) for many discussions and long-standing collaboration on this topic, Diogo Gomes and Alessandro Astolfi (KAUST) for their kind support, and J. Ricardo G. Mendonça (USP) for technical advice and a careful reading of the manuscript that greatly improved its presentation. This work was partially supported by the King Abdullah University of Science and Technology (KAUST), Saudi Arabia.